Adams Company makes and sells products with variable costs of $24 each. Adams incurs annual fixed costs of $321,280. The current sales price is $88.

Note: The requirements of this question are interdependent. For example, the $256,000 desired profit introduced in Requirement c also applies to subsequent requirements. Likewise, the $80 sales price introduced in Requirement d applies to the subsequent requirements.

Required

a. Determine the contribution margin per unit.

Complete this question by entering your answers in the tabs below.

Determine the break-even point in units and in dollars. (Do not round intermediate calculations. Round your final answers to the nearest dollar and round units up to the next whole unit.)

Break-even point in units
Break-even point in dollars

Prepare an income statement using the contribution margin format. (Do not round intermediate calculations. Round your final answers to nearest whole number.)

ADAMS COMPANY
Income Statement
Net income (loss)
Answer #1
a
Contribution margin per unit 64 =88-24
Fixed costs 321280
/ Contribution margin per unit 64
Break-even point in units 5020
Break-even point in dollars 441760 =5020*88
b
ADAMS COMPANY
Income Statement
Sales revenue 441760 =5020*88
Variable costs 120480 =5020*24
Contribution margin 321280
Fixed costs 321280
Net income(loss) 0